A heat exchanger wall consists of a copper plate 20 mm thick. The heat transfer

coefficients on the two sides of the plate are 2700 and 7000 W/m2K, corresponding to

fluid temperatures of 92 °C and 32 °C, respectively. Assuming that the thermal

conductivity of the wall is 375 W/mK,

a. Calculate the surface temperatures in °C, and

b. Calculate the overall heat transfer in W/m2K.


1
Expert's answer
2022-04-24T17:19:38-0400

"\frac qA=\frac{T_1-T_2}{\frac 1{h_1}+\frac Lk+\frac 1{h_2}}=105.9~\frac{kW}{m^2},"

"T_{w1}=T_1-\frac qA\frac 1{h_1}=52.8°C,"

"T_{w2}=T_2-\frac qA\frac 1{h_2}=47.1°C."